### v1

@@ Abstract
Dropout, a network operator, when enabled is likely to dramatically impact the performance of Flash-Attention, which in turn increases the end-to-end training time of Large-Language-Models (LLMs). The main contributor to such performance degradation is the Random Number Generation (RNG) phase that is traditionally fused into the Flash-Attention kernel. As RNG and Attention have the same hardware bottlenecks, RNG latency can hardly be hidden within the Attention kernel.
@@
We propose overlapping RNG with previous GEMM layers in the network to hide RNG runtime and improve end-to-end performance. RNG and GEMM have distinct resource requirements and hardware bottlenecks, so they can run in parallel without compromising each other’s performance. Our fine-grained performance model, cross-validated by silicon results, shows 1.14x speedup on one transformer block (including multi-head attention and feed-forward layers) for Llama2, and up to 1.23x speedup when varying workload sizes, on GH100 GPUs with FP8 precision. Further, we extend our theoretical model to different RNG implementations and hardware architectures, and discuss the widely applicable benefits for overlapping RNG with GEMM layers.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
@@
Large Language Models (LLMs) have become important targets for performance optimization due to their ever-increasing workload sizes and corresponding runtime demands. Full-scaled training for models like GPT OpenAI (2023) requires several months and thousands of GPUsII (2023). Attention dropout Srivastava et al. (2014)Zehui et al. (2019) is an optional technique that drops out elements after the Softmax operation in Attention. Dropout is applied in commonly used models such as Llama Face Touvron et al. (2023) because it can make the model focus on relevant features and improve training accuracy Schumacher Shastri  Xue et al. (2024). However, enabling dropout is costly, which doubles the processing time of the Attention layer with state-of-the-art implementations like Flash-Attention Dao et al. (2022)Dao (2023)Shah et al. (2024), and in turn increases the end-to-end training time by 1.3x to 1.7x, depending on the network parameters. Optimizing the runtime of dropout can significantly improve the performance of LLM training.
@@
The runtime of dropout is dominated by the Random Number Generator (RNG) Salmon et al. (2011), which generates random numbers to determine which elements within the intermediate matrix of the Attention layer to drop. The size of this matrix, and consequently the runtime of RNG, scales quadratically with the sequence length. As the industry trends towards ever larger sequence lengths, this exacerbates the RNG latency.
Traditionally, dropout (including RNG) is fused into the Flash-Attention kernel. The Attention layer is not limited by matrix-matrix multiplication (MMA) math but rather by non-compute bottlenecks, including the Register File read and write bandwidth, the Issue Stage, and Multi-Function Operations. Because RNG is also bottlenecked by these limiters, its latency is almost fully exposed when fused with Attention. Silicon measurements shows that only 10-20% of RNG runtime can be hidden with the Attention kernel for GPT-like workload sizes.
In this paper, we propose a method to overlap RNG with other matrix multiplication (GEMM) layers within the Transformer Block to hide RNG runtime and improve end-to-end performance. The stand-alone RNG kernel generates one bit for each element to decide whether it is dropped, and write them to the memory to be read and used later by the Attention kernel when performing the dropping.
@@
RNG and GEMM are ideal targets for overlapping because they do not depend on each other and have distinct resource requirements and hardware bottlenecks. GEMM is predominantly bounded by MMA math and L2 bandwidth, whereas RNG is mostly limited by the Issue Stage and ALU operations. Despite GEMM consuming nearly all the Register Files (RF) and Shared Memory (SMEM) within each Streaming Multiprocessor (SM), RNG uses minimal RF and SMEM, allowing GEMM to run in parallel with RNG while obtaining near-optimal performance.
Figure 1: Dropout, including RNG, is traditionally fused in Flash-Attention. Overlapping RNG with GEMM layers before Attention hides exposed RNG latency and results in speedup.
We built a fine-grained theoretical performance model for RNG and GEMM overlapping. This model analyzes hardware bottlenecks for individual kernels, fused kernels, and interference between overlapping kernels. It takes in GPU hardware configurations and the model parameters, and outputs the theoretical runtime for each kernel.
Our model is validated by silicon results running with FP8 precision NVIDIA (e) on H100 HBM3 80GB GPUs NVIDIA (f), showing only a 2% difference between theoretical and actual results. Results demonstrate great potential when overlapping RNG with the GEMM layers between the previous and the current Attention layer: 1.14x speedup for Llama2 Touvron et al. (2023), 1.13x for MoE (a trillion-parameter NVIDIA model prototype NVIDIA (h)) , and 1.06x for GPT3 Brown (2020). We also observed up to 1.23x speedup under varying network parameters within the range of common workloads.
We extend our analysis to different hardware designs to evaluate how changes in the chip’s computing power could impact results. We also explore the adaptability of the overlapping technique with different implementations of the RNG function. Furthermore, our overlapping methodology can be generalized to other layers in the network that meet three key criteria: no data dependencies (or resolvable through pipelining), no shared hardware bottlenecks, and no capacity conflicts. Possible targets for overlapping include the Communication layer, either with GEMM or even RNG, as well as the Matrix Transpose and Precision Conversion layers if required by the network.
@@
  • We proposed overlapping the key component of dropout, RNG, with GEMM kernels to hide RNG latency, leading to substantial speedup across multiple trending LLM architectures.
  • We developed a detailed theoretical performance model for assessing the benefits of overlapping RNG and GEMM, given the workload sizes of the network and the hardware specification. The accuracy of our model is cross-validated by results obtained from real silicon implementations.
  • We explored the broader implications of the overlapping mechanism for different RNG implementations, application on other targeting layers, and future hardware architecture designs.
2 Background
Figure 2: Network Architecture of one Transformer Block, the basic building block of LLM network. The four GEMM layers and the Attention layer dominate the runtime.
2.1 LLM Network Architecture
@@
In this work, we begin our analysis on single GPU without communication, and focus on the GEMM and Attention layers. We discuss multi-GPU scenarios in Section 5 and draw similar conclusions as in single GPU. With multi-GPU, different parallelism mechanisms NVIDIA (g) Shoeybi et al. (2019) are applied which evenly distribute workload onto different GPUs. This does not impact savings of overlapping since the ratio of each kernel’s runtime remains the same. In highly optimized implementations, communication layers are often overlapped with their producer layers to minimize latency overhead, therefore the end-to-end savings brought by overlapping still apply.
2.2 Dropout
@@
Figure 3: Dropout is applied to the intermediate results in Flash-Attention, after the Softmax operation.
@@
Dropout Srivastava et al. (2014)Wan et al. (2013) is a technique designed to prevent overfitting by randomly setting a small fraction of the elements within a network to zero. It has been proposed to regularize neuron networks in general. In the context of this paper, we refer specifically to Attention dropout Schumacher Shastri , where the dropout operation is applied to the intermediate outputs of the Attention layer following the Softmax operation, as depicted in Figure 3.
@@
While optional, dropout can substantially improve training accuracy: it prevents the model from relying heavily on certain features, thus making the model only focus on relevant features to prevent overfitting and underfitting Xue et al. (2024)Liu et al. (2023). Dropout is used in training widely adapted LLM networks such as Llama2 Face Touvron et al. (2023). However, its huge runtime slowdown has limited its application. We explore dropout optimizations to improve its practicality in future applications, facilitating broader implementation.
2.3 RNG Implementation: Philox
@@
The computational demand of dropout mostly comes from the Random Number Generator (RNG) used to determine which elements to zero out. Multiple possible RNG implementations exist, and our discussion centers on Philox Salmon et al. (2011) NumPy . Philox is a counter-based pseudorandom number generator (PRNG) that relies on wide multiplies and iterates over previous states to generate new ones. Philox is well-suited for GPU execution, as its operations can be efficiently parallelized. Implementations of Philox are available in NVIDIA’s cuRAND library NVIDIA (d) and TensorFlow TensorFlow . Our analysis primarily focuses on a seven-iteration implementation (Philox 7), though we also consider more resource-efficient versions (Philox 5 and 3) in Section 5.
3 Implementation
In this section, we discuss our experiments performed on silicon, as well as the fine-grained performance model which is backed up by silicon results. We first validate our performance assumptions that are essential for effective overlapping on silicon, then utilize the theoretical model to generalize our findings. Figure 4 shows the setup for baseline and overlap experiments.
Figure 4: Setup for baseline: dropout (including RNG) fused in Attention, and overlap: stand-alone RNG running in parallel with GEMM, with output consumed by Attention. D stands for the element dropping step.
3.1 Silicon Implementation
We tested a CUDA implementation of our proposal on NVIDIA H100 GPUs NVIDIA (f), the cutting-edge option for data-center training tasks, specifically using the GH100 HBM3 80GB variant. This GPU supports the latest High Bandwidth Memory (HBM) and provides ample compute resources and memory capacity for LLM training.
In terms of implementation, RNG and GEMM can either run as separate kernels or within the same kernel using warp specialization Li et al. (2023)NVIDIA (i)Kerr et al. . We opted for separate kernels to maintain a clear distinction between independent components in the network such as RNG and GEMM in this case. This approach allows other overlapping strategies to be implemented without deep knowledge of the kernels’ internal complexities, especially for the highly optimized GEMM kernels. Warp specialization enables a more fine-grained overlapping and we anticipate it to bring further, but limited, performance benefits.
Our silicon implementation uses production-ready, highly optimized GEMM, Attention, and dropout kernels. GEMM tile size is consistently set at 128x128x128, while RNG and Attention use a tile size of 64x128x128. The dropout kernel was modified to run in two scenarios: 1) fused within the Attention kernel, and 2) as a stand-alone RNG kernel storing bits representing random numbers in HBM for later use by the Attention kernel. We used the Philox 7 algorithm for RNG.
3.1.1 Performance Assumptions
To validate our assumptions about performance impacts, we conducted several tests:
  • GEMM Resource Allocation: Originally, the GEMM kernel utilizes all available Register Files and Shared Memory in each SM. We carved out 6% of the Registers and 7% of the Shared Memory for the RNG kernel, hypothesizing this would not adversely affect GEMM performance. Our silicon measurements confirmed only 0.5% average performance difference across various GEMM workload sizes.
3.1.2 Baseline and Overlap CUDA Implementation
The baseline implementation includes a stand-alone QKV_GEMM kernel and dropout (including RNG) fused into the Attention kernel, running serially on the same CUDA stream. Our optimized kernel fusion minimizes synchronization overhead and maximizes the overlap of RNG operations with Attention’s floating-point computations.
Both baseline and overlapping implementations use CUDAgraph NVIDIA (c) to minimize delays between kernel launches and optimize overall efficiency. The GPU is warmed up using multiple instances of the captured CUDA graphs before actual measurements are taken. We validated our CUDA implementation against CPU-generated result, ensuring correctness across various batch sizes, sequence lengths, and number of heads.
3.2 Theoretical Model
This subsection explains the construction of our fine-grained theoretical performance model, designed to evaluate the overlapping technique’s effectiveness based on hardware bottlenecks. We begin by discussing the hardware bottlenecks taken into account in our model, and then the derivation of baseline and overlapping runtime.
3.2.1 Hardware Limiters
@@
Table 1: Hardware Limiters.
Limiter Note
MMA Math (TFLOPS) Max compute intensity of Matmul, main limiter of GEMM.
L2 Bandwidth Read and write bandwidth between L2 and the SM cores.
@@
HBM Bandwidth Read and write bandwidth to the main memory.
RF Bandwidth Read and write bandwidth for the Register Files within each SM.
Instruction Issue Pipeline for issuing instructions.
ALU Pipe Pipeline for executing ALU instructions.
Multi-Function Units Pipe Pipeline for executing multi-function instructions.
FMA Pipe Pipeline for executing FMA instructions.
3.2.2 Modeled Layers

### v2

@@ Abstract.
Dropout, a network operator, when enabled is likely to dramatically impact the performance of Flash-Attention, which in turn increases the end-to-end training time of Large-Language-Models (LLMs). The main contributor to such performance degradation is the Random Number Generation (RNG) phase. The state-of-the-art optimization is to fuse RNG into the Flash-Attention kernel. However, while RNG and Attention do not compete on compute or memory resources, they are bounded on the same lower-level architecture bottlenecks. Fusion can hardly hide RNG latency within the Attention kernel.
@@
We propose overlapping RNG with previous GEMM layers in the network to hide RNG latency and improve end-to-end performance. RNG and GEMM have distinct resource requirements and hardware bottlenecks, so they can run together without compromising each other’s performance. We propose a fine-grained analytical performance model that analyzes low-level architecture resource utilization to evaluate RNG-GEMM overlapping performance benefits. This model, cross-validated by silicon results, shows 1.26x speedup for overlapping RNG and GEMM layers over a sequential implementation on one Transformer Block (one LLM layer including multi-head attention and feed-forward layers), and 1.22x over state-of-the-art fusion implementation, for Llama3 on GH100 GPUs with FP8 precision. Because the kernel patterns are regular, the findings of the shared bottlenecks, as well as the achievable performance benefits, can be generalized to different model architectures, software implementations and hardware configurations.
1. Introduction
@@
Large Language Models (LLMs) have become important targets for performance optimization due to their ever-increasing workload sizes and corresponding runtime demands. Full-scaled training for big models requires several months and thousands of GPUs (OpenAI, 2023)(II, 2023). Attention dropout (Srivastava et al., 2014)(Zehui et al., 2019) is an optional technique that drops out elements after the Softmax operation in Attention. Dropout is applied in commonly used models such as Llama (Face, [n. d.])(Touvron et al., 2023) because it can make the model focus on relevant features and improve training accuracy (Schumacher, [n. d.])(Shastri, [n. d.]) (Xue et al., 2024). However, enabling dropout is costly, which doubles the processing time of the Attention layer with state-of-the-art implementations like Flash-Attention (Dao et al., 2022)(Dao, 2023)(Shah et al., 2024), and in turn increases the end-to-end training time by 1.3x to 1.7x. Optimizing the runtime of dropout can significantly improve the performance of LLM training.
@@
The runtime of dropout is dominated by the Random Number Generator (RNG) (Salmon et al., 2011), which generates random numbers to determine which elements within the intermediate matrix of the Attention layer to drop. The size of this matrix, and consequently the runtime of RNG, scales quadratically with the sequence length. The industry trend towards ever larger sequence lengths further exacerbates the RNG latency.
Figure 1. Attention has a dependency on RNG and Dropout (“D”). State-of-the-art optimization is kernel fusion, which results in hardware bottleneck conflicts. We propose overlapping RNG with GEMM layers to achieve significant speedup.
A starter implementation is to run RNG sequentially before Attention, and store the random numbers to generate the masks for dropping out elements (“D”) in Flash-Attention, as shown in Figure 1. To improve performance, state-of-the-art solution fuses dropout (including RNG) into the Flash-Attention kernel. A common but misleading assumption is that Flash-Attention in LLM training is dominated by matrix-matrix multiplication (MMA) math and thus limited by compute capability. This assumption rules out the possibility that Attention and RNG share the same hardware bottlenecks, because RNG has no MMA operations. However, with a fine-grained analysis with more lower-level architecture bottlenecks on the kernel code, we observe that Attention is actually limited by non-compute hardware resources: the Register File bandwidth, and the Issue Stage bandwidth. Because RNG is also bottlenecked by the Issue Stage, its latency is almost fully exposed when fused with Attention. Silicon measurements show that only 10-20% of RNG runtime can be hidden within the Attention kernel.
@@
In this paper, we propose overlapping RNG with other matrix multiplication (GEMM) layers to hide RNG runtime and improve end-to-end performance. RNG and GEMM are ideal targets for overlapping because they do not depend on each other and have distinct resource requirements and hardware bottlenecks. GEMM is mostly bounded by MMA math and L2 bandwidth, whereas RNG mostly by the Issue Stage and ALU operations. While GEMM has a heavy consumption on the Register Files (RF) and Shared Memory (SMEM) within each Streaming Multiprocessor (SM), RNG uses minimal RF and SMEM, allowing GEMM to run in parallel with RNG while obtaining near-optimal performance.
We built a fine-grained analytical model to evaluate the performance for different kernel implementations: sequential execution, RNG-Attention fusion, and RNG-GEMM overlapping. This model analyzes hardware bottlenecks beyond compute and memory, and takes into account lower-level hardware configurations such as Register File, Issue Stage, and instruction execution pipelines. It takes in GPU hardware configurations and the model parameters, and outputs the runtime for kernels based on the workload’s primary bottleneck. Our model is validated by silicon results running with FP8 precision (NVIDIA, [n. d.]) on H100 HBM3 80GB GPUs (NVIDIA Corporation, 2022).
Results demonstrate great potential when overlapping RNG with the GEMM layers between the previous and the current Attention layer: for Llama3 (Grattafiori et al., 2024), overlapping achieves 1.26x speedup compared to baseline sequential kernels, and 1.22x when compared to state-of-the-art fusion methodology; similar speedup is obtained for GPT-4 (OpenAI, 2023). Since the identified bottlenecks for the analyzed kernels remain the same for regular LLM training workloads with dense, non-narrow matrix shapes, the performance benefits can be generalized.
Using the analytical model, we derive implications of RNG-GEMM overlapping technique under different hardware and software settings. As a general trend, future generations of GPUs advance in computing power (Tensor Cores), but are likely to have similar architecture designs for non-Tensor units such as Register File, Issue Stage and ALU pipelines. On these future GPUs, non-Tensor resource bottlenecks are even more severe, and overlapping RNG and GEMM is more favorable.
The analytical model is widely applicable with configurable inputs for model architecture (such as Sequence Length, Batch Size, Embedding Dimension), software implementations (such as tile size, kernel ordering), and hardware configurations (such as compute capability, memory bandwidth, Register File, Issue Stage and execution pipelines). Further, the model can be easily adapted to an universal overlapping/fusion scheduler evaluator that works for all operators.
@@
  • We proposed overlapping the key component of dropout - RNG - with GEMM kernels to hide RNG latency, leading to substantial speedup across multiple trending LLM models compared to state-of-the-art fusion technique.
  • We developed an analytical performance model based on fine-grained hardware bottlenecks beyond compute, communication and memory. This model accurately derives performance comparisons between sequential kernels, fused kernels and overlapping kernels. It is cross-validated by results obtained from real silicon implementations.
  • We explored the broader implications of the overlapping mechanism for various model architectures, future GPU hardware designs, and different software implementations.
2. Background
Figure 2. Network Architecture of one Transformer Block, the basic building block of LLM. The four GEMM and the Attention layers dominate the runtime, and the communication latency is typically hidden within the GEMM layers.
2.1. LLM Network Architecture
@@
In this work, we begin our analysis on single GPU without communication, and focus on the GEMM and Attention layers. We discuss multi-GPU scenarios in Section 7 and draw similar conclusions as in single GPU. With multi-GPU, different parallelism mechanisms (Shoeybi et al., 2019) are applied which evenly distribute workload onto different GPUs. This does not impact savings of overlapping since the ratio of each kernel’s runtime remains the same. In highly optimized implementations, communication layers are often overlapped with their producer layers to minimize latency overhead, therefore the end-to-end savings brought by overlapping still apply.
2.2. Dropout
@@
Figure 3. Dropout is applied to the intermediate results in Flash-Attention, after the Softmax operation.
@@
Dropout (Srivastava et al., 2014)(Wan et al., 2013) is a technique designed to prevent overfitting by randomly setting a small fraction of the elements within a network to zero. It has been proposed to regularize neuron networks in general. In the context of this paper, we refer specifically to Attention dropout (Schumacher, [n. d.])(Shastri, [n. d.]), where the dropout operation is applied to the intermediate outputs of the Attention layer following the Softmax operation, as depicted in Figure 3.
@@
While optional, dropout can substantially improve training accuracy: it prevents the model from relying heavily on certain features, thus making the model only focus on relevant features to prevent overfitting and underfitting (Xue et al., 2024)(Liu et al., 2023). Dropout is used in training widely adapted LLM networks such as Llama (Face, [n. d.])(Touvron et al., 2023). However, its huge runtime slowdown has limited its application. With our dropout optimizations, we can enable its usage in future applications.
2.3. RNG Implementation: Philox
@@
The computational demand of dropout mostly comes from the Random Number Generator (RNG) used to determine which elements to zero out. Multiple possible RNG implementations exist, and our discussion centers on Philox (Salmon et al., 2011) (NumPy, [n. d.]). Philox is a counter-based pseudorandom number generator (PRNG) that relies on wide multiplies and iterates over previous states to generate new ones. Philox is well-suited for GPU execution, as its operations can be efficiently parallelized. Implementations of Philox are available in NVIDIA’s cuRAND library (NVIDIA Corporation, [n. d.]) and TensorFlow (TensorFlow, [n. d.]). Our analysis primarily focuses on a seven-iteration implementation (Philox 7), though we also consider more resource-efficient versions (Philox 5 and 3) in Section 5.2.
3. Analytical Model Methodology
In this section, we explain the construction of our fine-grained analytical performance model. We discuss the hardware resources and limiters, the software implementations of LLM networks, and the derivation of runtime of individual kernels, kernel fusion and kernel overlapping.
3.1. Hardware Limiters
@@
Our model is at tile-level granularity because LLM implementations are usually tiled. For MatMul kernels, we first model one tile’s runtime, then derive the total runtime by multiplying with the total number of tiles. For non-MatMul operators such as stand-alone RNG, we directly model the entire kernel’s runtime based on the total number of operations and the primary bottlenecks.
Table 1. The hardware limiters included in our analytical model.
Limiter Hardware Resources Total Operations for Each Kernel
GEMM ($A_{m,k}B_{k,n}$) Attention ($Softmax(Q_{m,n}K_{m,n}^{T})V_{m,n}$) RNG ($m*m$ Dimension)
MMA Math Tensor Core #FLOP per second #MMA ops: $m*n*k$ Same as GEMM for two MatMuls: $m*n*m*2$ / (No MatMul)
@@
HBM Read BW Read bandwidth of the main memory #inputs: $m*k+n*k$ Same as GEMM for two Matmuls / (No memory reads)
L2 Read BW Read bandwidth between L2 cache and SM cores #$\#inputs*\#SMs$: $k(m*num\_n\_tiles+n*num\_m\_tiles)$ Same as GEMM for two Matmuls / (No memory reads)
RF Read BW Each SM’s Reg File read bandwidth #Types of instructions * #RF reads in each type
Issue Stage Pipeline for issuing instructions Number of instructions
ALU Pipe ALU instruction pipe Number of ALU instructions in kernel
MUFU Pipe MUlti-FUnction instruction pipe Number of multi-function unit instructions
FMA Pipe Fused Multiply-Add instruction pipe Number of FMA instructions
The hardware limiters that we chose to investigate in our model are listed in Table 1. These limiters are all configurable in order to model different hardware designs. For each row, runtime is calculated by the number of operations divided by the hardware resources available. Total operations for MatMuls are obtained from the theoretical number of mul-add operations based on matrix dimensions. To further improve model accuracy, with the access to production-ready, highly optimized kernel code, we obtain non-MatMul kernels’ number of operations by counting the number of instructions from the production kernel’s assembly code (SASS). These numbers can also be derived theoretically without the access to kernel assembly code.
  • MMA Math: Tensor Core TFLOPS obtained from GPU datasheet. GEMM and Attention kernels during LLM training perform dense matrix multiplications and are mostly bounded on MMA Math. Runtime is calculated by the number of MMA operations ($m*n*k$) divided by TFLOPS. RNG kernel does not contain MMA operations.
  • Main-memory Read BW: The memory (HBM in high-end datacenter GPUs) read bandwith in Bytes per second. Runtime is calculated by total number of memory reads divided by the memory bandwidth for GEMM and Attention layers’ matrix multiplications. RNG kernels only generate random numbers and there is no memory read.
  • L2 Read BW: Bandwidth for L2 cache reads. Cache reads for GEMM and Attention kernels are calculated as the size of each input Matrix multiplied by the number of SMs that access a certain element in the matrix. Since tiles are being computed on different SMs, each element needs to be read by multiple SMs. RNG does not read from memory nor does it have L2 reads.
  • RF Read BW: The register file read bandwidth inside each SM core. For all kernels, this is calculated by going through each type of the frequent instructions in the kernel, and obtaining the number of registers that each instruction type uses. For example, FMA (Fused Multiply-Add) instruction has three input registers. While GEMM typically uses a lot of registers due to its heavy matrix multiplication operations, this bottleneck is secondary compared to compute (MMA Math). However, Flash-Attention has softmax that heavily uses FMA and other operations with high demand of register file read bandwidth. The softmax demand on top of the $(QK^{T})V$ matrix multiplication makes register read bandwidth one of the most critical hardware bottlenecks for Flash-Attention.
  • Issue Stage: Typically one instruction is issued each cycle. Runtime is obtained by analyzing the number of total instructions in each kernel. RNG and Attention kernel (with softmax) both have a lot of instructions and are typically bounded by the issue stage.
  • ALU/MUFU/FMA Pipe: Number of ALU/MUFU/FMA instructions in the kernel divided by each function unit’s throughput. RNG kernel is heavily bounded on ALU pipe, while the Attention kernel has a lot of MUFU/FMA instructions.
We focus on read bandwidths because typically applications will have more reads than writes that go through any unit in the memory hierarchy. In practice, we also observe that when bandwidth becomes an issue, the real limiter is usually the read bandwidth but not the write.
3.2. Software Configurations
On the software implementation side, the model architecture we used is the vanilla Transformer architecture (Vaswani et al., 2017). Specifically, we model the QKV_GEMM, Flash-Attention (with softmax), PROJ_GEMM, FC1 and FC2 kernels, which are the main contributors to the runtime of a Transformer Block in LLM training. We do not model the communication layers because in practice, communication latency is shorter than the computation latency and they utilize completely different hardware resources. Communication latency can easily be hidden within the GEMM/Attention layers.
The configurable inputs to the model from the software implementation side include:
  • Batch Size
  • Input Sequence Length
  • Head Dimension
  • Embedding Dimension (= #Heads * Head Dimension)
  • Tile Size
3.3. Methodologies to model kernels